# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 2.11.5 Delivering current across the membrane
Following their arrival, the driving force perpendicular to the membrane’s voltage disappears, and the ions form a thin "hot spot" in the layer. The electric repulsion acts in the direction parallel with the membrane’s surface and leads to distributing the ions (decreasing the gradient by distributing the charge locally) around the channel’s exit. However, the ions have finite speed. The ions saturate the layer on the membrane’s surface with a time constant between ($\frac{10^{-8}m}{10^{-1}m/s}\ s$) at the beginning and $(\frac{10^{-8}m}{10^{-4}m/s}\ s)$ at the end of their arrival period (we assumed 10nm10\ nm average distance between ion channel exits on the membrane). We shall take the longer time, so that we can expect a time constant for the saturation current around the ion channel’s exit in the order of 0.1ms0.1\ ms. When charging up the membrane in an avalanche-like way, the ions must pass on average a distance of about 0.05mm0.05\ $m$ from its center to its farthest point, so we expect a 0.5ms0.5\ ms ($\frac{5*10^{-5}m}{10^{-1}m/s}\ s$) time until the membrane’s slow current charges up the membrane to its maximum potential. The created charge must flow out from the farthest point in the neuron membrane of size 0.1mm0.1\ $m$ in time of order at or below 1ms1\ ms ($\frac{10^{-4}m}{10^{-1}m/s}\ s$); see the length of the $\frac{dV}{dt}$ pulse measured at the beginning of the AIS [54], which time is prolonged up to 10ms10\ ms by the neuronal $RC$ circuit; the ions are slow when the voltage on the AIS is low, see Eq. (25). Assuming those distances and speeds, including the potential-assisted speed of the slow current, we are on a time scale matching the available observations.
@@ 2.11.6 Synaptic gating
The slow current also explains the phenomenon of synaptic gating. The synaptic charge-up current flows into the membrane. It causes transient changes [55, 56] in the membrane’s voltage, providing direct evidence that the membrane is not always equipotential. The ions on the membrane’s surface can propagate at a finite speed. Given that the ions in the axonal arbor need to enter the membrane against the actual membrane potential, the membrane potential acts as a valve. The potential stops the ion inflow to the membrane for the period while the membrane’s voltage is above the threshold: it effectively inhibits further inflow through all synapses. This behavior naturally explains the absolute refractory period. After the membrane’s voltage drops below the threshold value, the ions can enter the membrane again (and the current intensity increases as the membrane’s potential decreases), but they need time to reach the AIS later when in the meantime the membrane’ voltage proceeded toward its hyperpolarized state; so they seem to appear dozens of microseconds later at the AIS, explaining the relative refractory period (as a delayed expansion of the absolute refractory period, with variable current intensity). Of course, the length of the relative refractory period sensitively depends on the conduction velocity [51].
@@ 2.11.7 Voltage sensing
"Voltage sensing by ion channels is the key event enabling the generation and propagation of electrical activity in excitable cells."[57] The mechanism by which voltage-gating of channels works remains a mystery; one of the worst consequences of the HH theory is that it separates the potential from ions and their current. It is not easy to investigate it experimentally: "the structural basis of voltage gating is uncertain because the resting state exists only at deeply negative membrane potentials" [58]. Usually, a "sliding helix" (structural) model is assumed.
@@
Under certain conditions, an ion channel can be opened in only one direction and only for a limited period, and this allows the membrane to become semipermeable. We imagine an ion channel as a simple hole (a cylinder) between the high and low-concentration segments with a cap on its top (on the side of the low-concentration segment). At the spots where the ion channels are located, the ions cannot penetrate the membrane until the cap is removed/lifted (the channel gets open). Unlike the original Maxwell demon, our demon does not have information in advance about which particle should be transmitted: it is passive in selecting the particle. (Passive here means that no biologically produced local energy is used: the electric potential energy from the voltage difference across the membrane moves the ions to the other side of the membrane.)
@@
Our voltage-controlled ion channel gets opened and closed due to purely electrostatic reasons. It works as a two-plate simple nano-scale electrometer (of type quadrant, Lindermann, Hoffman, and Wulf), similar to the ones used to measure the small electrical potential between charged elements (e.g., plates or fine quartz fibers). Given that the membrane and the cap in their resting state are isolators, no electrical repulsion is evoked between them, and the adhesion sticks them firmly together, representing a permanent force. The van der Waals force is inversely proportional to the squared distance between the dipoles in the cap and the membrane, respectively, and is linearly proportional to the perimeter of the channel.
@@
However, when a slow ion current flows into the surface layer in the proximity of the cap, charges appear in the layer proximal to the membrane; the membrane and the cap get covered by a fragile electrical skin. The charge on the cap is proportional to the surface of the cap and similarly inversely proportional to the squared distance between the cap and the membrane. A local voltage gradient is generated by the local gradient of the slow ion current, and the force acting on the cap is proportional to the product of the voltage gradient and the area of the cap. Given that the cap is slightly elevated, the repulsion force may have a component in the direction of lifting the cap. Since the van der Waals force has a fixed size, when the electrical repulsion exceeds it at a critical voltage gradient value, the channel opens.
@@
The gate remains open as long as the local charge distribution enables it. The cap is connected to the membrane only at one point, so it cannot fly away, and also cannot close again until the charge on the surface is present. In the absence of charge, the cap undergoes random movement, and the short-distance van der Waals force may eventually reposition the cap on the membrane, thereby closing the channel. The voltage-sensing electrometer opens the channel, and the lack of charge on the surface enables it to close, but the closing is not immediate (the mass of the cap is by orders of magnitude larger than the mass of an ion that can pass the channel). The fluctuation of the voltage gradient due to the slow current in the layer in the proximity of the membrane near the ion channel’s exit opens, closes, and re-opens the channel in an apparently stochastic way (actually, as the repulsion of charges due to the fluctuating current on the cap and the membrane regulates it), as observed.
@@
When one cap is removed, the rushed-in ions in the proximity of the channel’s exit suddenly increase the local potential (produce fast transient changes [56]) proximal to the spot centered at the exit in the layer on the membrane’s surface. The surface outside the spot remains at a lower potential, so the ions in the layer start moving toward other channel exits, delivering potential to those channel exits. Given that they are voltage-controlled, they get open, and the process continues in an avalanche-like way [59]. The avalanche, as explained, needs a sufficiently large voltage gradient, which several synaptic inputs can trigger if they sum up appropriately. Alternatively, a single spike with sufficiently steep front slope [60] can be sufficient; providing a simple way of synchronizing neuronal assemblies (and proving that not the voltage, but the voltage gradient, single of summed, controls the operation).
@@
The operation of the ion channel alone cannot explain why the channel closes after a given number of ions have passed the channel; that number is not (entirely) random. The local behavior of the membrane’s surface layers regulates the number of ions that pass through. The segments are no longer mechanically separated when the cap is removed. The charged ions are enabled to rush into the lower concentration segment. They experience an enormous accelerating gradient; that, comparable to that of electrostatic particle accelerators, "snorts" the ions from the high-concentration side into the low-concentration side and causes a process "like a flee hopping in a breeze" [25]. Consequently, "transport efficiency of ion channels is $10^{5}$ times greater than the fastest rate of transport mediated by any known carrier protein" [45]. Having gates (caps) is needed only for the synchronized operation of the channels: opening and closing work also in the absence of caps, as discussed by [27].
@@
The snorted ions "hop" into the layer from another layer. In the beginning, with their voltage-accelerated speed, it could take less than $\frac{5*10^{-9}m}{10^{3}m/s}\ s$ to pass the channel (simulation [61] uses a psecpsec representative time interval), in the end, they may slow down to the voltage-assisted level as the potential gradually decreases (which is still $\frac{5*10^{-9}m}{10^{-1}m/s}\ s$), so we can omit that time when calculating the charged layer formation (as discussed in Appendix A, the estimated speed is about $10^{5}\ m/s$). Due to the enormous speed difference between the accelerated and assisted speeds, the passage is practically instant. On the high-concentration segment, only the ions in the layer in the immediate proximity of the entrance can feel the accelerating potential and move with the potential-accelerated speed. The after-diffusion with the potential-assisted speed from the next neighboring layer in the high potential segment is by orders of magnitude slower than the passage through the hole with the potential-accelerated speed.
@@
Here, the effect of the finite resources explicitly appears: about $10^{3}$ ions are transferred per channel. These ions are snorted from one layer in the high-concentration segment into another layer in the low-concentration layer. The driving force gradually decreases because ions leave the first layer and appear in the second layer. The potential-assisted speed to replace the leaving ions in the first segment from the bulk as well as diffusing out from the second segment without appropriate driving forces, this process is by orders of magnitude slower; therefore, we can approximate the process as one where a gradually decreasing accelerating force drives the ions. In a very short period, the potential difference between the layers disappears (see Fig. 7) that prevents further ions from entering the formerly high-concentration layer: the process of transferring ions through the channel closes the door behind the needed amount of ions. Now the gradient diminished, and the van der Waals force can close the channel again.
  • the potential generated across the membrane is entirely neglected
  • the ions have no driving force and the hypothesized protein carriers are too slow [45]
  • the considered van der Waals force is too weak to be noticed by the ions
  • the assumed force by the ’cation-attractive negative ends’ destabilizes the ion path: as the deviation from the central path increases, so increases the deviating driving force
  • even if the weak van der Waals force would work for a single ion, the next ion would be rejected by the strong Coulomb-force due to the first ion

### v3

@@ 4.10.3 Delivering current across the membrane
Following their arrival, the driving force perpendicular to the membrane’s voltage disappears, and the ions form a thin "hot spot" in the layer. The electrical repulsion acts in the direction parallel with the membrane’s surface and leads to distributing the ions (decreasing the gradient by distributing the charge locally) around the channel’s exit. However, the ions have a finite speed. The ions saturate the layer on the membrane’s surface with a time constant between ($\frac{10^{-8}m}{10^{-1}m/s}\ s$) at the beginning and $(\frac{10^{-8}m}{10^{-4}m/s}\ s)$ at the end of their arrival period (we assumed 10nm10\ nm average distance between ion channel exits on the membrane). We shall take the longer time, so that we can expect a time constant for the saturation current around the ion channel’s exit in the order of 0.1ms0.1\ ms. When charging up the membrane in an avalanche-like way, the ions must pass on average a distance of about 0.05mm0.05\ $m$ from its center to its farthest point, so we expect a 0.5ms0.5\ ms ($\frac{5*10^{-5}m}{10^{-1}m/s}\ s$) time until the membrane’s slow current charges up the membrane to its maximum potential. The created charge must flow out from the farthest point in the neuron membrane of size 0.1mm0.1\ $m$ in time of order at or below 1ms1\ ms ($\frac{10^{-4}m}{10^{-1}m/s}\ s$); see the length of the $\frac{dV}{dt}$ pulse measured at the beginning of the AIS [53], which time is prolonged up to 10ms10\ ms by the neuronal $RC$ circuit; the ions are slow when the voltage on the AIS is low, see Eq. (28). Assuming those distances and speeds, including the potential-assisted speed of the slow current, we are on a time scale matching the available observations.
4.10.4 Ion selectivity
The mechanism of channel passing can also contribute to explaining ion selectivity. "The normal selectivity cannot be explained by pore size, because $Na^{+}$ is smaller than $K^{+}$ [41]". The fact that ion selectivity changes with the membrane potential is observed (under the name of changed specific permeability), but it remains unexplained. "For example, the negative resting potential of nerve cells is largely determined by a class of $K^{+}$ channels that are 100-fold more permeable to $K^{+}$ than to $Na^{+}$. In contrast, during the action potential a class of $Na^{+}$ channels is activated that are 10- to 20-fold more permeable to $Na^{+}$ than to $K^{+}$. Thus, a key to the great versatility of neuronal signaling is the regulated activation of different classes of ion channels, each of which is selective for specific ions." [17], page 101. However, the question of why the permeability changes and how the channels are activated remains open. It is remarkable that an ion-specific channel is permeable also for other ions, only different transmission intensity is observed.
The commonly used picture of the operation of selectivity filters is surely wrong. The assumed mechanical operation of the pores is too slow: the assumed structural change needs $10^{-8}\ s$ and the ions passage time is about $10^{-11}\ s$ (furthermore, it must be repeated about $10^{3}$ times per passage). If a wrong ion is caught, it must be transported back to its departure side, through the right ions (against their repulsion), and the right ions must retry. Neither forward nor backward movement has an appropriate driving force in that picture. Furthermore, the transport capacity is not sufficient for the operation.
In our model, as Eq. (6) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamic forces are equal. The $E_{thermal}^{C_{k}}(d)$ "electrical field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions differ. While a resulting force is exerted on one ion and moves it to the other segment, the other remains resting, or even moves in the opposite direction. Furthermore, the rush-in of ions greatly increases the $Na^{+}$ concentration in the dynamic layer, naturally increasing the number of $Na^{+}$ ions passing through the channel. That increase alone creates the illusion of the increased $Na^{+}$ permeability. Of course, the ions moving across the segments change the field contributions, indirectly affecting the balanced state of the other ions. The permanently open ion channels participate passively in the process, and their apparent permeability changes with the applied fields. To provide channel penetrability data is meaningful only if ion concentrations and membrane thickness data are also available. The driving forces of the ions differ, and they produce, depending on the actual concentrations and local potential, the selectivity of the channels.
4.10.5 Ion pumps
The role of ion pumps must also be revisited. "These pumps differ from ion channels in two important details. First, whereas open ion channels have a continuous water-filled pathway through which ions flow unimpeded from one side of the membrane to the other, each time a pump moves an ion, or a group of a few ions, across the membrane, it must undergo a series of conformational changes. As a result, the rate of ion flow through pumps is 100 to 100,000 times slower than through channels. Second, pumps that maintain ion gradients use energy, often in the form of ATP, to transport ions against their electrical and chemical gradients. Such ion movements are termed active transport." [17], page 101.
It is not clear what a force can "transport ions against their electrical and chemical gradients". It is undoubtedly "not on the ground that there is any ‘new force’ or whatnot" [7] present: the ions, independently of whether they are in an inanimate or living environment, can be moved only by electrical or thermodynamic forces, instead of some magic protein mechanism (ion pumps). The resultant potential (the sum of the electrical and thermodynamic potential for the individual ions) moves the ions through the channels in the membrane. In a typical mammalian cell, see Table 1, the sum of the electrical and thermodynamic potentials is $32\ [mV]\ K^{+}$ ions and $4\ [mV]\ Na^{+}$ ions. The same values for squids are $33\ [mV]\ K^{+}$ and $13\ [mV]\ Na^{+}$. Correspondingly, in the resting state, the resulting force moves sodium ions out of the cell and potassium ions into the cell (for more numerical examples see section 6.1); the same way as the "Na-K pump" is supposed to do by using some hypothesized protein mechanisms. In the classical model, no mechanism is provided for how the ion pumps gain energy from ATP and convert it into forward force. In our model, the potential gradient near the membrane, the ATP, by hydrolysis, generates ions, which move to the "plates" of the condenser, thus increasing its potential energy. That energy is prepared "globally" (not at the proximity where it is used) and continuously (during the full life cycle of neuronal operation), mainly in the recovery period. As discussed in section 5.2, in the transient state, to restore the condenser’s voltage, those ions must be collected.
In our model, the ion pumps are ion channels which work in stationary rather than transient conditions. In the resting state, since the resultant forces exerted on the different ions differ, they permanently pass through the membrane. As described, the passage changes the ionic gradients on the two sides of the membrane and generates a slight "leakage current" and slightly fluctuating voltage. In the resting state, the electrolyte layers have sufficient time for after-diffusion, allowing a graded ionic gradient to form near the membrane. (The gradient changes in a fraction of a nanometer, so it is challenging to measure it.) With reference to Fig. 7, the graded gradient generates an electrical field on the ion channels, which is by orders of magnitude lower than before rush-in. Conversely, in the transient mode, the capped ion channels enable the creation of ungraded (step-like) gradients. Consequently, the ion speed, furthermore, the charge and mass transport, are by order of magnitude lower in the resting state than in the transient state, as observed.
As Figure 6 in [43] displays, the ratio of $K+$ and $Na+$ changes sharply during the transient state. Theoretically, they assumed that the neuron “pumps 3 $Na^{+}$ ions out of the cell and two potassium ions in”; experimentally, they show in their Fig. 6 that the ratio changes between 0.01 and 7.5. According to our theory, the magnitude of the current (and so the exchange ratio) depends on the resulting potentials, so the ratio must change as the membrane potential changes during AP. The classical theory cannot explain this behavior and leads to exciting conclusions: "Furthermore, we analyzed energy properties of each ion channel and found that, under the two circumstances, power synchronization of ion channels and energy utilization ratio have significant differences. This is particularly true of the energy utilization ratio, which can rise to above 100% during subthreshold activity." [43] Our model says that (due to $Na^{+}$ rush-in) both the concentration and local membrane potential drastically change during the transient state. As long as the resulting potential is above $\approx 33\ mV$, the driving force acting on potassium ions is positive, i.e., the pump will move both ions out of the cell.
@@ 4.10.6 Synaptic gating
The slow current also accounts for synaptic gating. The synaptic charge-up current flows into the membrane. It causes transient changes [54, 55] in the membrane’s voltage, providing direct evidence that the membrane is not always equipotential. The ions on the membrane’s surface can propagate at a finite speed. Given that the ions in the axonal arbor need to enter the membrane against the actual membrane potential, the membrane potential acts as a valve. The potential stops the ion inflow to the membrane for the period while the membrane’s voltage is above the threshold: it effectively inhibits further inflow through all synapses. This behavior naturally explains the absolute refractory period. After the membrane’s voltage drops below the threshold value, the ions can enter the membrane again (and the current intensity increases as the membrane’s potential decreases), but they reach the AIS later when in the meantime the membrane’ voltage proceeded toward its hyperpolarized state; so they seem to appear dozens of microseconds later at the AIS, explaining the relative refractory period (as a delayed expansion of the absolute refractory period, with variable current intensity). Of course, the length of the relative refractory period sensitively depends on the conduction velocity [48].
@@ 4.10.7 Voltage sensing
"Voltage sensing by ion channels is the key event enabling the generation and propagation of electrical activity in excitable cells."[56] The mechanism by which voltage-gating of channels works remained a mystery; one of the worst consequences of the HH theory is that it separates the potential from ions and their current. It is not easy to investigate it experimentally: "the structural basis of voltage gating is uncertain because the resting state exists only at deeply negative membrane potentials" [57]. Usually, a "sliding helix" (structural) model is assumed.
@@
Under certain conditions, an ion channel can be opened in only one direction and only for a limited period, and this allows the membrane to become semipermeable. We imagine an ion channel as a simple hole (a cylinder) between the high and low-concentration segments with a cap on its top (on the side of the low-concentration segment). At the spots where the ion channels are located, the ions cannot penetrate the membrane until the cap is removed/lifted (the channel gets open). Unlike the original Maxwell demon, our demon does not have information in advance about which particle should be transmitted: it is passive in selecting the particle. (Passive here means that no biologically-produced local energy is used: the electrical potential energy from the voltage difference across the membrane moves the ions to the other side of the membrane.)
@@
In our model, a voltage-controlled ion channel gets opened and closed due to purely electrostatic reasons. It works as a two-plate simple nano-scale electrometer (of type quadrant, Lindermann, Hoffman, and Wulf), similar to the ones used to measure the small electrical potential between charged elements (e.g., plates or fine quartz fibers). Given that the membrane and the cap in their resting state are isolators, no electrical repulsion evokes between them, and the adhesion sticks them firmly together, representing a permanent force. The van der Waals force is inversely proportional to the squared distance between the dipoles in the cap and the membrane, respectively, and is linearly proportional to the perimeter of the channel.
@@
However, when a slow ion current flows into the surface layer in the proximity of the cap, charges appear in the layer proximal to the membrane; the membrane and the cap get covered by a fragile electrical skin. The charge on the cap is proportional to the area of the cap and similarly inversely proportional to the squared distance between the cap and the membrane. The local gradient of the slow ion current generates a local voltage gradient, and the force acting on the cap is proportional to the product of the voltage gradient and the area of the cap. Given that the cap is slightly elevated, the repulsion force may have a component in the direction of lifting the cap. Since the van der Waals force has a fixed size, when the electrical repulsion exceeds it at a critical voltage gradient value, the channel opens.
4.10.8 Gated ion channels in action
@@
The gate remains open as long as the local charge distribution enables it. The cap is connected to the membrane only at one point, so it cannot fly away, and also cannot close again as long as the charge on the surface is present. In the absence of charge, the cap undergoes random motion, and the short-range van der Waals force may eventually reposition it on the membrane, thereby closing the channel. The voltage-sensing electrometer opens the channel, and the lack of charge on the surface enables it to close, but the closure is not immediate (the cap’s mass is by orders of magnitude larger than the mass of an ion that can pass through the channel). The fluctuation of the voltage gradient due to the slow current in the layer in the proximity of the membrane near the ion channel’s exit opens, closes, and re-opens the channel in an apparently stochastic way (actually, as the repulsion of charges due to the fluctuating current on the cap and the membrane regulates it), as observed.
@@
When one cap is removed, the rushed-in ions in the proximity of the channel’s exit suddenly increase the local potential (produce fast transient changes [55]) proximal to the spot centered at the exit in the layer on the membrane’s surface. The surface outside the spot remains at a lower potential, so the ions in the layer start moving toward other channel exits, delivering potential to those channel exits. Given that those channels are voltage-controlled, they get open, and the process continues in an avalanche-like way [58]. The avalanche, as explained, needs a sufficiently large voltage gradient, which several synaptic inputs can trigger if they sum up appropriately. Alternatively, a single spike with sufficiently steep front slope [59] can be sufficient, providing a simple way to synchronize neuronal assemblies (and proving that not the voltage, but the voltage gradient, single or summed, controls the operation).
4.10.9 Gating layers
@@
The operation of the ion channel alone cannot explain why the channel closes after a given number of ions have passed through the channel; that number is not (entirely) random. The local behavior of the membrane’s surface layers regulates the number of ions that pass through. The segments are no longer mechanically separated when the cap is removed. The charged ions are enabled to rush into the lower concentration segment. They experience an enormous accelerating gradient; that, comparable to that of an electrostatic particle accelerator, the neuron "snorts" the ions from the high-concentration side into the low-concentration side and causes a process "like a flee hopping in a breeze" [34]. Consequently, "transport efficiency of ion channels is $10^{5}$ times greater than the fastest rate of transport mediated by any known carrier protein" [41]. Having gates (caps) is needed only for the synchronized operation of the channels: opening and closing work also in the absence of caps, as discussed by [35].
@@
The snorted ions "hop" into the layer from another layer. At the beginning, with their voltage-accelerated speed, it could take less than $\frac{5*10^{-9}m}{10^{3}m/s}\ s$ to pass the channel (simulation [60] uses a psecpsec representative time interval), in the end, they may slow down to the voltage-assisted level as the potential gradually decreases (which is still $\frac{5*10^{-9}m}{10^{-1}m/s}\ s$), so we can omit that time when calculating the charged layer formation (as discussed in Appendix B, the estimated speed is about $10^{4}~to~10^{5}\ m/s$). Due to the enormous speed difference between the accelerated and assisted speeds, the passage is practically instant. On the high-concentration segment, only the ions in the layer in the immediate vicinity of the entrance can feel the accelerating potential and move with the potential-accelerated speed. The after-diffusion with the potential-assisted speed from the next neighboring layer in the high potential segment is by orders of magnitude slower than the passage through the hole with the potential-accelerated speed.
@@
Here, the effect of the finite resources explicitly appears: about $10^{3}$ ions are transferred per channel. These ions are snorted from one layer in the high-concentration segment into another layer in the low-concentration layer. The driving force gradually decreases because ions leave the first layer and appear in the second layer. The potential-assisted speed to replace the leaving ions in the first segment from the bulk, as well as to diffuse out from the second segment without appropriate driving forces, is by orders of magnitude slower; therefore, we can approximate the process as one in which a gradually decreasing accelerating force drives the ions. In a very short period, the potential difference between the layers disappears (see Fig. 7) that prevents further ions from entering the formerly high-concentration layer: the process of transferring ions through the channel closes the door behind the needed amount of ions. Now the gradient diminished, and the van der Waals force can close the channel again.
4.11 Fallacies about biological electricity
4.11.1 Ion transport
In balanced states, no transport occurs, so the transport force cancels, and the mobility has no role (when speaking about mobility provided by protein mechanisms, one should derive the corresponding contribution to Eq. (7)). The counterforce adapts to the situation. That force may be mechanical: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions on the opposite side, and the membrane mechanically provides the needed counterforce. This way, no resultant force acts on the ions; the two gradients persist. When ions can move freely between segments, they continue to move until they produce a concentration gradient (a thermodynamic force) for the given ion that counterbalances the electrical gradient (the electrical force), at which point transport stops. A transport force is needed to reach the Stokes-Einstein speed (see Eq. (30)) in a viscous fluid.
We must not forget that initially, the Nernst-Planck equation described a transfer process (i.e., a dynamic equation), assuming the same speed for mass and charge transfers. We used it only to describe the balanced state (i.e., a static equation at zero speed) and added the term describing external (such as mechanical constraint) force that may be involved; it affects the process, but it does not belong to either of the respective fields. Commonly known mechanical constraints appear in a non-conducting layer called the membrane, where that counterforce prevents ions from penetrating the membrane layers. A membrane is a perfect isolator, i.e., no charge carriers exist between its two surfaces. (There may be ion channels that deliver ions built into the membrane, but it is a different subject.) The charge-up results from charge separation, a mechanism distinct from polarization, as typically described in most biology textbooks. In the form we derived, the equation correctly describes ion transfer, a current conveying ions with the speed controlled by the actual gradients, instead of assuming an ion current traveling with the apparent speed of electrical current in solids.
4.11.2 Modelling
It was a colossal mistake to forget that biological currents consist of ions, each with its electrostatic force. It led to fallacies, among others, including delayed and leakage currents (see Section 4.5.4) and heat absorption (see Section 7.2). Their finite (Stokes-Einstein) speed, see section 4.8, due to the ions’ mass and the different conduction mechanism, explains the spatiotemporal behavior of living matter [61] (contrasted to Newton’s ’instant interaction’).
It was another colossal mistake to introduce equivalent circuits with constant-value voltage generators. It forces one to assume that the conductances of the biological objects (membrane, synapses, AIS) change without any reason, and prevents understanding how the competition between thermodynamic and electrical processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 4.8; this way attributing the change of an electrical entity to the biological material. This assumption neglects that a driving force is needed to move ions (denies Newton’s laws of motion). Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
4.11.3 Measuring
The classic textbook [40], appendix A, correctly warned (especially when working with concepts of an electrical theory of neuronal operation): "An intuitive grasp of concepts of electricity and electrical circuits is helpful for understanding some of the basic theory in cellular neurophysiology. …Under ideal circumstances, the physical act of measuring a neurophysiological event would have no effect on the electrical signal of interest. Unfortunately, this is seldom the case in neurophysiology."
As discussed in section 6.1, physiology does not realize that, in resting and transient states, different physical processes take place and correspondingly, different circuit models shall be used. In the resting state, the low-intensity ion current is "shunted" by the ion channels in the membrane’s wall; moreover, the "parallel RCRC" model works sufficiently well. Since the current is slow, it is not simply distributed, but only the "overflow" reaches the AIS. However, as the current intensity exceeds the ion-conductance capacity of the "resting ion channels", the "transient ion channels" play an increasingly important role. Above the voltage threshold, the serial $RC$ circuit is the correct model. Until the neuron’s voltage reaches its threshold, a "parallel" and a "serial" resistor work simultaneously, this way implementing a mixed model (for the correct mathematical handling, see Eq. (44)). Measuring the "input resistance" of neurons actually measures, in this sense, a mixture of the "input" and "output" resistance corresponding to the two models, given that the current is distributed between the one through the membrane and the other through the AIS; plus the measuring electrode represents one more drain.
At the beginning, it was clearly told [50], but forgotten: "Since it is quite generally believed that the depolarization of a nerve fiber membrane, during excitation and propagation, involves an increased permeability to ions. There have been many attempts to detect and measure this change as an increase in the electrical conductivity. … In these cases, the measuring current was also the stimulating current, and it was not possible to analyze the changes satisfactorily." Moreover, during the passage of an impulse, a "charge wave" (in another terminology: a "pressure wave") passes along the axon and the changed number of charge carriers in that cross-section, see Eq.(20), changes the current that creates the illusion that the conductance at the measured cross-section changed. In our terminology, the stabilizing feedback voltage provides an external force, and it also contributes an external current. The feedback amplifier’s electron current (converted to ion current) is provided externally and is added to the closed system, without including the measurement device into the current budget of the closed system. That interpretation of a closed system denies the fundamental symmetries of science (among them, the conservation of charge, mass, momentum, and energy). Moreover, it significantly contributes to the claim that "laws of science cannot describe life".
4.11.4 The ’outlaw’ biology
All conservation laws are correct and valid, only "the construction is different" [7] and the "test method" is wrong. Not mentioning that under some conditions, the "system under test" also "produces" charge carriers by opening ion channels. The "construction" needs revisiting concepts of electricity for living matter instead of projecting concepts valid for inanimate matter only, to living matter. That wrong "grasp of concepts of electricity" suggests that Ohm’s Law is not valid for living systems. The correct statement is that Ohm’s Law is not valid for arbitrarily chosen currents and voltages. Without that revisiting, by saying that living systems show non-Ohmic behavior, biology claims that it applies good laws inappropriately to a good system. The consequences are far-reaching: physiology uses magically changing conductances in circuits with fixed-voltage batteries, that blocks not only understanding the thermodynamic and mechanical effects (and introducing the corresponding theories [10, 62, 8, 63, 64, 35, 65, 66, 18, 67, 9, 68]), but even consistently describing the electrical phenomena. We are not speaking against using the concept of conductance, but rather against deriving its value incorrectly by including foreign current components, which obscures the actual electrical behavior of biological systems. In addition, by introducing current feedback (which, by definition, compensates for any gradient in the system), physiology applies an opposite-phase control unit against the neuron’s native control unit and concludes that, after compensation, there is no gradient in the system; in other words, that life exists without needing a driving force.
Neuroscience observes the thermodynamic and mechanical changes (for a comprehensive review, see [62]) but cannot connect them to the electrical phenomena. Lacking the required physical background, biology claims that physical laws cannot describe processes in living matter, without providing a plausible and self-consistent theoretical model.